What is the boiling point elevation of 45.6 g of NaCl dissolved in 650 g of water?


1
Expert's answer
2021-04-07T11:31:13-0400

deltaT = Kbm

where:

       deltaT is the change in the boiling point of the solvent,

        Kb is the molal boiling point elevation constant (for water Kb=0.512oC⋅m-1)

        m is the molal concentration of the solute in the solution.

Cm = n/m

n = m/M

M (NaCl) = 58.44 g/mol

n (NaCl) = 45.6 / 58.44 = 0.8 mol

Cm (NaCl) = 0.8 / 0.65 = 1.2 m

deltaT = 0.512⋅1.2 = 0.6oC